OpenStudy (anonymous):

The sine of a 30 degree angle is equal to the cosine of a 60 degree angle. true or false?

OpenStudy (zzr0ck3r):

\(\sin(\frac{\pi}{2}+\theta)=\cos(\theta)\)

OpenStudy (trisaba):

true

OpenStudy (anonymous):

thank you!!

OpenStudy (trisaba):

no problem